OpenStudy (anonymous):

A store sells batteries in packets of 6 or 10. It has a stock of 25 packets containing a total of 186 batteries. How many of these are 10-battery packets?

OpenStudy (bahrom7893):

So suppose you have x packets of 6 and y packets of 10. The total number of batteries would be: 6*x + 10*y = 186

OpenStudy (bahrom7893):

In total, you have 25 packets, so the sum of all the packets containing 6 and 10 batteries is: x+y=25

OpenStudy (kc_kennylau):

And we don't usually write the asterisk (*) if it's not necessary, so it'd be 6x+10y=186

OpenStudy (bahrom7893):

Solve the system: 6x+10y=186 x+y=25 y will be the number of 10-battery packets